# GALE: Leveraging Heterogeneous Systems for Efficient Unstructured Mesh Data Analysis

### v1

@@ 1.1 Simplicial Complex
A $k$-simplex (or simplex of dimension $k$) is defined as the convex hull of $k+1$ linearly independent points in the Euclidean space. A 00-simplex corresponds to a point, a $1$-simplex is an edge, a $2$-simplex forms a triangle, and a $3$-simplex represents a tetrahedron. Given a $k$-simplex $\sigma$, the convex hull of a nonempty subset of size $m+1$ of the $k+1$ points (i.e., $m<k$) that defines an $m$-simplex $\tau$ is called an $m$-face of $\sigma$, and $\sigma$ is said to be a coface of $\tau$. The set of cofaces of a simplex $\sigma$ forms the star of $\sigma$. In general, 00-faces are also known as vertices, $1$-faces are called edges, and $(n-1)$-faces are called facets. For instance, a 3-simplex (or tetrahedron) contains four 0-faces (vertices), six 1-faces (edges), and four 2-faces (facets/triangles).
@@ 1.2 Topological Relations
There are three major groups of topological relations that describe the connectivity of the simplices in a simplicial complex $\Sigma$. The boundary relation maps a simplex to its faces, the coboundary relation maps a simplex to its cofaces, and the adjacency relation maps a simplex to adjacent simplices of the same dimension. Suppose that two simplices $\sigma$ and $\tau$ are in $\Sigma$, and $\sigma$ is a face of $\tau$, we say that $\sigma$ is on the boundary of $\tau$, and similarly, $\tau$ is on the coboundary of $\sigma$. Two $k$-simplices $\tau_{1}$ and $\tau_{2}$ are adjacent if and only if they share a common $(k-1)$-simplex $\sigma$, and two vertices are adjacent if they are on the same edge.
In this paper, we focus on the topological relations inside a tetrahedral mesh and use capital letters to indicate whether the relation involves a vertex ($V$), edge ($E$), triangle ($F$), or tetrahedron ($T$). Each topological relation is denoted as a pair of letters, e.g., the $FE$ relation represents the edges on the boundary of a triangle. Table 1 shows the topological relations existing in a tetrahedral mesh, which includes six boundary relations, six coboundary relations, and four adjacency relations.
@@
Figure 1 illustrates a toy example demonstrating the use of topological relations to compute a derived value from a scalar field. Specifically, for each vertex in the tetrahedral mesh, a new value is obtained by summing the scalar values of its adjacent vertices. Figure 1(a) depicts a scalar field defined over a simplicial mesh consisting of two tetrahedra sharing a common triangular face. In Figure 1(b), the $VV$ relation for the vertex $v_{0}$ is computed as $VV(v_{0}):[v_{1},v_{2},v_{3}]$. Finally, Figure 1(c) shows the $VV$ relation is used to compute the derived value for $v_{0}$ by summing the scalar values of its neighboring vertices.
@@ 1.3 Task Parallel Approach for Mesh Processing
Traditional visualization algorithms process data in two sequential phases. First, the necessary connectivity information is computed, and then it is used to derive the desired results [44]. For example, Figure 1(b) illustrates the data computation phase, where the $VV$ relation is computed, while Figure 1(c) depicts the data consumption phase, where each vertex’s derived scalar value is computed as the sum of the scalar values of its adjacent vertices. Figure 2(a) further illustrates how these two phases are executed in a sequential implementation: the $VV$ relations are first computed globally for each vertex, and only then is this information utilized to produce the final results.
@@
The idea of task-parallel approaches is to run these two phases in parallel by assigning them to different groups of threads. This task-parallel approach is realized with a producer-consumer design [42], where a producer thread computes the data connectivity, while a consumer thread uses the connectivity data to compute the desired information. In the example shown in Figure 2(b), this corresponds to producer and consumer threads working on different vertices in parallel.
@@ 2 Related Work
In this section, we review existing work on data structures for unstructured meshes and parallel algorithms for topological mesh processing. We categorize data structures into three groups: global data structures that process the entire mesh at once (Section 2.1), local data structures that handle individual chunks of the mesh (Section 2.2), and GPU data structures specifically designed to leverage GPU acceleration (Section 2.3). While some data structures prioritize geometric information [45], they do not support the computation of topological relations and are therefore excluded from our discussion.
@@ 3 A Task-parallel Approach for Hybrid CPU-GPU Computing
This section first outlines the key challenges of the proposed model in Section 3.1, followed by an overview of the data structure in Section 3.2. We then introduce our concrete implementation, the GPU-Aided Localized data structurE (GALE), with a detailed discussion of its input and initialization phase in Section 3.3. Subsequent sections describe the information flow between consumers (Section 3.4) and producers (Section 3.5). Additional implementation details can be found in supplemental materials.
@@ 3.2 Overview of GALE
To enable efficient mesh processing, we propose GALE (GPU-Aided Localized data structurE), a task-parallel data structure that explicitly addresses the challenges outlined in Section 3.1 while leveraging both CPU and GPU resources.
Figure 3 provides an overview of GALE’s threading model and communication pipeline. GALE is designed around three key components: consumer threads, a leader producer, and worker producers, each fulfilling a distinct role in the execution pipeline.
Consumer threads are CPU threads responsible for running the data analysis algorithm. These threads pause execution whenever a topological relation is required. GALE’s first design choice is to distinguish between boundary relations, which are computed directly by the consumer without GPU acceleration, and coboundary or adjacency relations, which are delegated to producer threads. The behavior of consumer threads is described in detail in Section 3.4.
Leader producer is a CPU thread that facilitates communication between the CPU and GPU, aggregating requests from consumers, launching the computation kernel, and synchronizing data transfers. This middle-layer design choice reduces synchronization overhead and improves task coordination. Consumer-to-leader communication is handled via FIFO queues, and leader-to-consumer communication is realized through semaphores, as detailed in Section 3.4, while leader-to-GPU interactions are described in Section 3.5.
@@ 3.3 Data Encoding
Figure 4 shows an example of a tetrahedra mesh composed of six vertices and three tetrahedra, along with the corresponding $V$, $T$, and $S$ arrays. The segmentation defined by $S$ indicates the mesh is divided into two segments. Vertices $v_{0}$, $v_{1}$, and $v_{2}$ belong to segment $S_{1}$, while the remaining vertices belong to segment $S_{2}$. GALE maps each simplex $\sigma$ to a segment $S_{i}$. A simplex $\sigma$ is internal to the segment $S_{i}$ iff the vertex $v$ of $\sigma$ with the lowest index also belongs to $S_{i}$ and is external to all other intersecting segments. For example, in Figure 4, tetrahedron $t_{1}$ is internal to segment $S_{1}$ because $S[v_{0}]=1$, but external to segment $S_{2}$ because it contains $v_{4}$, where $S[v_{4}]=2$. While GALE supports any vertex-based subdivision, our experimental evaluation employs the PR octree technique [38]. This method segments the mesh by defining a maximum number of vertices per leaf node, ensuring general applicability to any spatially embedded mesh [15].
@@
The initialization phase prepares the data structure after reading the input data but before algorithm execution. Figure 5 illustrates the additional data encoded after initialization for the input mesh in Figure 4. The initialization phase involves enumerating mesh edges and triangles, a process performed on the CPU. Enumeration is carried out by iterating through the tetrahedra in each mesh segment. For each tetrahedron, all unique combinations of two vertices (edges) and three vertices (triangles) are stored in sorted indexed lists ($E$ and $F$) based on vertex indices. $T_{ex}$ encodes the list of external tetrahedra for each segment, which is used for efficiently computing coboundary relations (see LABEL:sec:gpu-computation). The data structure also maintains an interval array $I$ for each type of simplex, where $[I[S_{k}-1],\ I[S_{k}])$ defines the range of indices of all internal simplices for segment $S_{k}\ (k>0)$. This information is used to quickly look up the segment containing a given simplex (see Section 3.4).
@@ 3.4 Consumer Threads
If the relation is a boundary relation, the consumer computes it only for $\sigma$. GALE optimizes this process via a few lookup operations in its initialized data structures. For example, to compute the $FE$ relation for triangle $f_{5}$ in Figure 4: (1) Retrieve three boundary vertices from the triangle list $F$: $v_{0}$, $v_{3}$, and $v_{4}$. (2) Locate edge indices for $(v_{0},v_{3})$, $(v_{0},v_{4})$, and $(v_{3},v_{4})$ through the edge list $E$ with limiting the search range by the edge interval array $I_{E}$. For instance, since $v_{0}$ belongs to the first segment, the index of $(v_{0},v_{3})$ is searched from the position $I_{E}[0]=0$ to $I_{E}[1]=9$, while the index of $(v_{3},v_{4})$ is searched between $I_{E}[1]=9$ and $I_{E}[2]=12$. (3) The final results for $FE(f_{5})$ consist of edge indices $e_{2}$, $e_{6}$, and $e_{9}$.
If the relation is a coboundary or adjacency relation, the consumer delegates its computation to the GPU via the flow outlined in Figure 3. In this case, the consumer request is not limited to the simplex $\sigma$ but involves the entire segment $S_{k}$. Consumer requests are organized into dedicated queues, one for each coboundary or adjacency relation. The consumer acquires the lock to the queue and pushes its request containing information about the mesh segment and the consumer thread index. Section 3.5 elaborates on the queue system, clarifying how it addresses the challenge [C1] between the consumer and leader producer.
@@ 3.5 Queue System and Worker Producers
A dedicated leader producer thread is spawned for each queue to establish a one-to-one correspondence. When a queue contains incoming requests, the corresponding leader producer processes them in batches and launches the corresponding computation kernel on the GPU. As described in Section 1.3, a key factor in improving the performance of task-parallel approaches is to ensure the connectivity data is prepared in advance for consumer threads. To this end, the workload sent to the GPU by the leader producer includes not only the currently requested segments but also subsequent segments for proactive precomputation. The total number of segments computed by the GPU kernel is defined as $Q_{r}\cdot n_{b}\cdot t_{b}/t_{s}$, where $Q_{r}$ is the number of consumer requests popped from the queue, while $n_{b}$, $t_{b}$ and $t_{s}$ are all user-defined parameters. Specifically, $n_{b}$ specifies the number of GPU blocks used to process a request, $t_{b}$ denotes the number of GPU threads for each block, and $t_{s}$ indicates the number of threads allocated for processing a segment of the mesh. To avoid the overhead of repeated memory allocation and deallocation associated with the dynamic computation of topological relations, the system pre-allocates sufficient GPU memory for each consumer thread, based on the number of segments to be precomputed (i.e., $n_{b}\cdot t_{b}/t_{s}$).
Figure 6 shows an example of the leader producer preparing a GPU kernel launch with $n_{b}=2$. In this example, the leader producer collects three segments from consumer requests in the queue (Figure 6a), which are $S_{4}$, $S_{71}$, and $S_{102}$. Based on the input parameters, the system computes 4 segments per request, distributed across 2 GPU blocks. The leader producer allocates memory space in the relation array for requested segments and their subsequent 3 segments (Figure 6b), then distributes the workload across GPU blocks (Figure 6c). Once the kernel finishes execution and transfers the connectivity data back to the main memory, the leader producer maps the simplices in the relation array from their local indices to global indices, integrating the data into the format used by the algorithm.
@@ 4 Experimental evaluation
To evaluate performance across different computational scenarios, we selected three topological data analysis (TDA) algorithms from TTK for benchmarking. We chose TDA algorithms because they are typically more memory-intensive and resource-demanding than standard geometric mesh analysis algorithms. Section 4.1 details the experimental setup, while Section 4.2 compares our approach with state-of-the-art topological data structures. Finally, Section 4.3 analyzes the efficiency of our design, demonstrating how it successfully addresses the challenges outlined in Section 3.1.
@@ 4.1 Experimental Setup
All experiments were conducted on a computer cluster node equipped with two 20-core Intel® Xeon Gold 6148 CPUs, 64 GB of RAM, and an NVIDIA® Tesla V100 GPU. We use six tetrahedral meshes listed in Table 2. The Fish and Hole datasets are unstructured tetrahedral meshes, while the remaining four tetrahedral meshes, Engine, Foot, Asteroid, and Stent, are generated by removing null values from regular volume datasets followed by tetrahedralization. All the datasets have been partitioned using the PR octree [38].
@@
DiscreteGradient. This algorithm [44] computes a discrete gradient field [16] based on the input scalar field. Intuitively, a discrete vector field resembles a collection of arrows connecting a $k$-simplex of mesh $\Sigma$ to an incident $(k+1)$-simplex in such a way that each simplex is either the head or tail of at most one arrow and a critical simplex is neither the head nor the tail of any arrow. The algorithm first iterates all vertices in the mesh, and for each vertex, it adds all $k$-simplices ($k>0$) in the lower star of the vertex into a list. If the list is not empty, the algorithm finds the 1-simplex (i.e., edge) to pair for the vertex, 2-simplices (i.e., triangles) for the remaining 1-simplices, and so forth. This process requires eight topological relations: $VE$, $VF$, $VT$, $EV$, $FV$, $FE$, $TV$, and $TF$ relations. GALE generates three queues and three leader producers for coboundary relations $VE$, $VF$, and $VT$ to handle the workload. Although still embarrassingly parallel, this algorithm demands significantly more connectivity information, posing a heavier computational load on the GPU.
@@ 4.2.1 Analysis of performance
For this algorithm, GALE consistently demonstrates the fastest data structure across all datasets, with ACTOPO requiring more time than GALE and TTK Explicit Triangulation being the slowest one to finish the algorithm, as shown in Figure 7. Overall, GALE achieves a 1.6×1.6$\times$1.6 × speedup over ACTOPO and a 4.7×4.7$\times$4.7 × speedup over TTK Explicit Triangulation. Regarding memory usage, GALE uses only about 5% more memory than ACTOPO on average and saves approximately 52% memory compared to TTK Explicit Triangulation. GALE and ACTOPO exhibit better memory performance than Explicit Triangulation due to their localized data structures, though their memory management strategies differ. The buffering system of ACTOPO allocates memory based on a percentage of the total number of segments, whereas GALE precomputes and stores a fixed number of segments. As a result, ACTOPO may require more memory for larger datasets, such as Hole and Stent. GALE is the fastest data structure because it maximizes the number of CPU threads dedicated to consumer tasks, further optimizing performance compared to ACTOPO.
@@
Figure 8 presents the time and memory comparisons. GALE achieves an overall 1.2×1.2$\times$1.2 × speedup over ACTOPO and 5.1×5.1$\times$5.1 × speedup over TTK Explicit Triangulation. In terms of memory efficiency, ACTOPO remains the most memory-efficient structure. GALE, with its expanded set of topological relations, uses approximately 20% more memory than ACTOPO. However, TTK Explicit Triangulation, which computes and stores all global relation arrays in memory, increases memory usage by an average of 70% compared to GALE.
@@
The algorithm lacks parallelism, and thus all data structures utilize only a single consumer during the execution. However, the task-parallel approach enables ACTOPO and GALE to benefit from some level of parallelism, as the remaining threads can still function as worker producers. Given that the algorithm is output-sensitive, performance results exhibit significant variability, and Figure 9 presents the time and memory usage results. GALE outperforms ACTOPO, primarily due to faster topological relation computation on the GPU, achieving an average speedup of 2.7×2.7$\times$2.7 ×. For most datasets, GALE even surpasses TTK Explicit Triangulation, delivering an average speedup of 2.1×2.1$\times$2.1 ×. The exception is the Stent dataset, where GALE is slower than TTK Explicit Triangulation but still remains comparable in performance.

### v3

@@ 1.1 Simplicial Complex
[…]
@@ 1.2 Topological Relations
[…]
In this paper, we focus on the topological relations inside a tetrahedral mesh and use capital letters to indicate whether the relation involves a vertex ($V$), edge ($E$), triangle ($F$), or tetrahedron ($T$). Each topological relation is denoted as a pair of letters, e.g., the $FE$ relation represents the edges on the boundary of a triangle. Table˜1 shows the topological relations existing in a tetrahedral mesh, which includes six boundary relations, six coboundary relations, and four adjacency relations.
@@
Figure˜1 illustrates a toy example demonstrating the use of topological relations to compute a derived value from a scalar field. Specifically, for each vertex in the tetrahedral mesh, a new value is obtained by summing the scalar values of its adjacent vertices. Figure˜1(a) depicts a scalar field defined over a simplicial mesh consisting of two tetrahedra sharing a common triangular face. In Figure˜1(b), the $VV$ relation for the vertex $v_{0}$ is computed as $VV(v_{0}):[v_{1},v_{2},v_{3}]$. Finally, Figure˜1(c) shows the $VV$ relation is used to compute the derived value for $v_{0}$ by summing the scalar values of its neighboring vertices.
@@ 1.3 Task Parallel Approach for Mesh Processing
Traditional visualization algorithms process data in two sequential phases. First, the necessary connectivity information is computed, and then it is used to derive the desired results [44]. For example, Figure˜1(b) illustrates the data computation phase, where the $VV$ relation is computed, while Figure˜1(c) depicts the data consumption phase, where each vertex’s derived scalar value is computed as the sum of the scalar values of its adjacent vertices. Figure˜2(a) further illustrates how these two phases are executed in a sequential implementation: the $VV$ relations are first computed globally for each vertex, and only then is this information utilized to produce the final results.
@@
The idea of task-parallel approaches is to run these two phases in parallel by assigning them to different groups of threads. This task-parallel approach is realized with a producer-consumer design [42], where a producer thread computes the data connectivity, while a consumer thread uses the connectivity data to compute the desired information. In the example shown in Figure˜2(b), this corresponds to producer and consumer threads working on different vertices in parallel.
@@ 2 Related Work
In this section, we review existing work on data structures for unstructured meshes and parallel algorithms for topological mesh processing. We categorize data structures into three groups: global data structures that process the entire mesh at once (Section˜2.1), local data structures that handle individual chunks of the mesh (Section˜2.2), and GPU data structures specifically designed to leverage GPU acceleration (Section˜2.3). While some data structures prioritize geometric information [45], they do not support the computation of topological relations and are therefore excluded from our discussion.
@@ 3 A Task-parallel Approach for Hybrid CPU-GPU Computing
This section first outlines the key challenges of the proposed model in Section˜3.1, followed by an overview of the data structure in Section˜3.2. We then introduce our concrete implementation, the GPU-Aided Localized data structurE (GALE), with a detailed discussion of its input and initialization phase in Section˜3.3. Subsequent sections describe the information flow between consumers (Section˜3.4) and producers (Section˜3.5). Additional implementation details can be found in supplemental materials.
@@ 3.2 Overview of GALE
To enable efficient mesh processing, we propose GALE (GPU-Aided Localized data structurE), a task-parallel data structure that explicitly addresses the challenges outlined in Section˜3.1 while leveraging both CPU and GPU resources.
Figure˜3 provides an overview of GALE’s threading model and communication pipeline. GALE is designed around three key components: consumer threads, a leader producer, and worker producers, each fulfilling a distinct role in the execution pipeline.
Consumer threads are CPU threads responsible for running the data analysis algorithm. These threads pause execution whenever a topological relation is required. GALE’s first design choice is to distinguish between boundary relations, which are computed directly by the consumer without GPU acceleration, and coboundary or adjacency relations, which are delegated to producer threads. The behavior of consumer threads is described in detail in Section˜3.4.
Leader producer is a CPU thread that facilitates communication between the CPU and GPU, aggregating requests from consumers, launching the computation kernel, and synchronizing data transfers. This middle-layer design choice reduces synchronization overhead and improves task coordination. Consumer-to-leader communication is handled via FIFO queues, and leader-to-consumer communication is realized through semaphores, as detailed in Section˜3.4, while leader-to-GPU interactions are described in Section˜3.5.
@@ 3.3 Data Encoding
Figure˜4 shows an example of a tetrahedra mesh composed of six vertices and three tetrahedra, along with the corresponding $V$, $T$, and $S$ arrays. The segmentation defined by $S$ indicates the mesh is divided into two segments. Vertices $v_{0}$, $v_{1}$, and $v_{2}$ belong to segment $S_{1}$, while the remaining vertices belong to segment $S_{2}$. GALE maps each simplex $\sigma$ to a segment $S_{i}$. A simplex $\sigma$ is internal to the segment $S_{i}$ iff the vertex $v$ of $\sigma$ with the lowest index also belongs to $S_{i}$ and is external to all other intersecting segments. For example, in Figure˜4, tetrahedron $t_{1}$ is internal to segment $S_{1}$ because $S[v_{0}]=1$, but external to segment $S_{2}$ because it contains $v_{4}$, where $S[v_{4}]=2$. While GALE supports any vertex-based subdivision, our experimental evaluation employs the PR octree technique [38]. This method segments the mesh by defining a maximum number of vertices per leaf node, ensuring general applicability to any spatially embedded mesh [15].
@@
The initialization phase prepares the data structure after reading the input data but before algorithm execution. Figure˜5 illustrates the additional data encoded after initialization for the input mesh in Figure˜4. The initialization phase involves enumerating mesh edges and triangles, a process performed on the CPU. Enumeration is carried out by iterating through the tetrahedra in each mesh segment. For each tetrahedron, all unique combinations of two vertices (edges) and three vertices (triangles) are stored in sorted indexed lists ($E$ and $F$) based on vertex indices. $T_{ex}$ encodes the list of external tetrahedra for each segment, which is used for efficiently computing coboundary relations (see LABEL:sec:gpu-computation). The data structure also maintains an interval array $I$ for each type of simplex, where $[I[S_{k}-1],\ I[S_{k}])$ defines the range of indices of all internal simplices for segment $S_{k}\ (k>0)$. This information is used to quickly look up the segment containing a given simplex (see Section˜3.4).
@@ 3.4 Consumer Threads
If the relation is a boundary relation, the consumer computes it only for $\sigma$. GALE optimizes this process via a few lookup operations in its initialized data structures. For example, to compute the $FE$ relation for triangle $f_{5}$ in Figure˜4: (1) Retrieve three boundary vertices from the triangle list $F$: $v_{0}$, $v_{3}$, and $v_{4}$. (2) Locate edge indices for $(v_{0},v_{3})$, $(v_{0},v_{4})$, and $(v_{3},v_{4})$ through the edge list $E$ with limiting the search range by the edge interval array $I_{E}$. For instance, since $v_{0}$ belongs to the first segment, the index of $(v_{0},v_{3})$ is searched from the position $I_{E}[0]=0$ to $I_{E}[1]=9$, while the index of $(v_{3},v_{4})$ is searched between $I_{E}[1]=9$ and $I_{E}[2]=12$. (3) The final results for $FE(f_{5})$ consist of edge indices $e_{2}$, $e_{6}$, and $e_{9}$.
If the relation is a coboundary or adjacency relation, the consumer delegates its computation to the GPU via the flow outlined in Figure˜3. In this case, the consumer request is not limited to the simplex $\sigma$ but involves the entire segment $S_{k}$. Consumer requests are organized into dedicated queues, one for each coboundary or adjacency relation. The consumer acquires the lock to the queue and pushes its request containing information about the mesh segment and the consumer thread index. Section˜3.5 elaborates on the queue system, clarifying how it addresses the challenge [C1] between the consumer and leader producer.
@@ 3.5 Queue System and Worker Producers
A dedicated leader producer thread is spawned for each queue to establish a one-to-one correspondence. When a queue contains incoming requests, the corresponding leader producer processes them in batches and launches the corresponding computation kernel on the GPU. As described in Section˜1.3, a key factor in improving the performance of task-parallel approaches is to ensure the connectivity data is prepared in advance for consumer threads. To this end, the workload sent to the GPU by the leader producer includes not only the currently requested segments but also subsequent segments for proactive precomputation. The total number of segments computed by the GPU kernel is defined as $Q_{r}\cdot n_{b}\cdot t_{b}/t_{s}$, where $Q_{r}$ is the number of consumer requests popped from the queue, while $n_{b}$, $t_{b}$ and $t_{s}$ are all user-defined parameters. Specifically, $n_{b}$ specifies the number of GPU blocks used to process a request, $t_{b}$ denotes the number of GPU threads for each block, and $t_{s}$ indicates the number of threads allocated for processing a segment of the mesh. To avoid the overhead of repeated memory allocation and deallocation associated with the dynamic computation of topological relations, the system pre-allocates sufficient GPU memory for each consumer thread, based on the number of segments to be precomputed (i.e., $n_{b}\cdot t_{b}/t_{s}$).
Figure˜6 shows an example of the leader producer preparing a GPU kernel launch with $n_{b}=2$. In this example, the leader producer collects three segments from consumer requests in the queue (Figure˜6a), which are $S_{4}$, $S_{71}$, and $S_{102}$. Based on the input parameters, the system computes 4 segments per request, distributed across 2 GPU blocks. The leader producer allocates memory space in the relation array for requested segments and their subsequent 3 segments (Figure˜6b), then distributes the workload across GPU blocks (Figure˜6c). Once the kernel finishes execution and transfers the connectivity data back to the main memory, the leader producer maps the simplices in the relation array from their local indices to global indices, integrating the data into the format used by the algorithm.
@@ 4 Experimental evaluation
To evaluate performance across different computational scenarios, we selected three topological data analysis (TDA) algorithms from TTK for benchmarking. We chose TDA algorithms because they are typically more memory-intensive and resource-demanding than standard geometric mesh analysis algorithms. Section˜4.1 details the experimental setup, while Section˜4.2 compares our approach with state-of-the-art topological data structures. Finally, Section˜4.3 analyzes the efficiency of our design, demonstrating how it successfully addresses the challenges outlined in Section˜3.1.
@@ 4.1 Experimental Setup
All experiments were conducted on a computer cluster node equipped with two 20-core Intel® Xeon Gold 6148 CPUs, 64 GB of RAM, and an NVIDIA® Tesla V100 GPU. We use six tetrahedral meshes listed in Table˜2. The Fish and Hole datasets are unstructured tetrahedral meshes, while the remaining four tetrahedral meshes, Engine, Foot, Asteroid, and Stent, are generated by removing null values from regular volume datasets followed by tetrahedralization. All the datasets have been partitioned using the PR octree [38].
@@
[…]
@@ 4.2.1 Analysis of performance
For this algorithm, GALE consistently demonstrates the fastest data structure across all datasets, with ACTOPO requiring more time than GALE and TTK Explicit Triangulation being the slowest one to finish the algorithm, as shown in Figure˜7. Overall, GALE achieves a 1.6×1.6$\times$1.6 × speedup over ACTOPO and a 4.7×4.7$\times$4.7 × speedup over TTK Explicit Triangulation. Regarding memory usage, GALE uses only about 5% more memory than ACTOPO on average and saves approximately 52% memory compared to TTK Explicit Triangulation. GALE and ACTOPO exhibit better memory performance than Explicit Triangulation due to their localized data structures, though their memory management strategies differ. The buffering system of ACTOPO allocates memory based on a percentage of the total number of segments, whereas GALE precomputes and stores a fixed number of segments. As a result, ACTOPO may require more memory for larger datasets, such as Hole and Stent. GALE is the fastest data structure because it maximizes the number of CPU threads dedicated to consumer tasks, further optimizing performance compared to ACTOPO.
@@
Figure˜8 presents the time and memory comparisons. GALE achieves an overall 1.2×1.2$\times$1.2 × speedup over ACTOPO and 5.1×5.1$\times$5.1 × speedup over TTK Explicit Triangulation. In terms of memory efficiency, ACTOPO remains the most memory-efficient structure. GALE, with its expanded set of topological relations, uses approximately 20% more memory than ACTOPO. However, TTK Explicit Triangulation, which computes and stores all global relation arrays in memory, increases memory usage by an average of 70% compared to GALE.
@@
The algorithm lacks parallelism, and thus all data structures utilize only a single consumer during the execution. However, the task-parallel approach enables ACTOPO and GALE to benefit from some level of parallelism, as the remaining threads can still function as worker producers. Given that the algorithm is output-sensitive, performance results exhibit significant variability, and Figure˜9 presents the time and memory usage results. GALE outperforms ACTOPO, primarily due to faster topological relation computation on the GPU, achieving an average speedup of 2.7×2.7$\times$2.7 ×. For most datasets, GALE even surpasses TTK Explicit Triangulation, delivering an average speedup of 2.1×2.1$\times$2.1 ×. The exception is the Stent dataset, where GALE is slower than TTK Explicit Triangulation but still remains comparable in performance.